OpenStudy (anonymous):

What is the arc length when theta=pi/3 and the radius is 5cm? A.10pi/3 cm B.16/3cm C.pi/3cm D.5pi/3cm

OpenStudy (fibonaccichick666):

so what is the formula for arc length?

OpenStudy (anonymous):

I'm not sure

OpenStudy (anonymous):

Assuming \(\theta\) is the central angle of a sector given in radians, you can set up a ratio equation: \[\frac{\text{arc length}}{\text{central angle}}=\frac{\text{circumference}}{\text{one revolution}}~~\implies~~\frac{L}{\theta}=\frac{2\pi r}{2\pi}\] which says \(L=r\theta\), where \(r\) is the radius of the given circle and \(L\) is the length of the arc subtended by the central angle \(\theta\).